# Singular bifurcations in a slow-fast modified Leslie-Gower model with Holling type II functional response, weak Allee effect and a generalist predator

### v1

@@ 6.3. Exit Chart
Refer to caption $\epsilon_{3}$$\epsilon_{3}$$v_{3}$$v_{3}$$r_{3}$$r_{3}$$\ell_{3}$$\ell_{3}$$\mathcal{M}_{3}$$\Delta_{3}^{\textup{in}}$$\Delta_{3}^{\textup{out}}$
Figure 11. Dynamics of (35) in the chart $K_{3}=\{\bar{u}=1\}$ for $A-M+AM-\frac{1}{Q}<0$ on the left and $A-M+AM-\frac{1}{Q}>0$ on the right. The flow in blue occurs in the $(\epsilon_{3},v_{3})$-plane, while the green one in the $(r_{3},v_{3})$-plane. The black orbits are a sketch of a sample of orbits flowing from $\Delta_{3}^{\textup{in}}$ and arriving at $\Delta_{3}^{\textup{out}}$. The red surface is the center manifold $\mathcal{P}_{3}$. Notice that since $\ell_{3}$ is simply the critical manifold $\mathcal{M}_{0}^{1}$ written in the local coordinates of this chart, the orbits on $\mathcal{P}_{3}$ can be identified with the slow manifolds that perturb from $\mathcal{M}_{0}^{1}$. Although we do not show $\Delta_{c}^{\textup{in}}$, as it is an exponentially thin strip, it is evident from the analysis presented that some of the orbits near $\mathcal{P}_{3}$ arrive at $\Delta_{3}^{\textup{out}}$.
@@ 6.4. Bottom chart
$u=r_{4}u_{4},\quad v=-r_{4},\quad\varepsilon=r_{4}\epsilon_{4},$ (58)
@@
$\begin{split}\frac{\textup{d}r_{4}}{\textup{d}t}&=r_{4}\epsilon_{4}(Q+u_{1})(A^{2}M+\mathcal{O}(r_{4})),\\ \frac{\textup{d}u_{4}}{\textup{d}t}&=-AMQ(1+(A+M-AM)u_{4})u_{4}-A^{2}M(Q+u_{4})u_{4}\epsilon_{4}+\mathcal{O}(r_{4}),\\ \frac{\textup{d}\epsilon_{4}}{\textup{d}t}&=-\epsilon_{4}^{2}(Q+u_{4})(A^{2}M+\mathcal{O}(r_{4})).\end{split}$ (59)
@@
Refer to caption $\epsilon_{1}$$u_{4}$$r_{4}$
Figure 12. Dynamics of (59) near the origin. In contrast to the flow of (35), the center manifold in the plane $(v_{4},\epsilon_{4})$ is not unique.
6.5. Blown-up dynamics at $T_{\mathcal{C}}$, relaxation osillations, and transitory canard
@@
At the singular level, i.e. for $\varepsilon=0$, the map $\Pi_{0\to 2}$ is described by the concatenation of the corresponding orbits for $\bar{r}=0$ in the blow-up analysis performed in Sections 6.1-6.4, refer in particular Figures 9, 10 and 11 and the analysis leading to them, together with a blow-down. For the map $\Pi_{1\to 2}$ one just needs to be careful to have the appropriate parameters such that the slow flow along $\mathcal{M}_{0}^{1,r}$ is directed towards $T_{\mathcal{C}}$ as in Figure 7 and that the parameter $Q$ is chosen so that there is a maximal canard at the fold point $P$. This is achieved for parameters exponentially close to $Q=Q_{H}$. Since we have already shown in Section 5 that the singular Hopf bifurcation at $P$ for $Q=Q_{H}$ may be supercritical, the appropriate choice of parameters guarantee that such a maximal canard is stable for $\varepsilon>0$ sufficiently small. ∎
@@
For $A-M+AM<\frac{1}{Q}$, one can show that trajectories that cross $\Sigma_{0}$ transversally converge to the origin. Since this does not lead to oscillatory behavior, we do not detail this case anymore. It is nevertheless interesting to observe that, in this case, the degenerate point $T_{\mathcal{C}}$ is a saddle but eventually attracts all orbits.
@@
Consider system (2) and let $C=-AMQ$ and $A-M+AM>\frac{1}{Q}$.
@@
Figure 15. Simulations of (2) with $S=\varepsilon=0.05$, $A=\frac{1}{2}$, $M=-\frac{1}{10}$, and $C=-AMQ$. The equilibrium point $E_{1}$ is located close to the fold point $P$. The parameter $Q$ is chosen as $Q=Q_{H}-\delta$ where $\delta$ is shown at the top of each figure. Notice, indeed, the canard explosion occurring for the chosen parameters. We highlight that some of the canard cycles pass through the degenerate point $T_{\mathcal{C}}$.
@@ 8. Numerical analysis and validation
For our simulations, we take the following setup of values of the parameters: $A=\frac{1}{2}$, $M=-\frac{1}{10}$, $\varepsilon=\frac{1}{20}$, while $(C,Q)$ is going to be bifurcation parameters. In Figure 17, we present a codimension $2$ bifurcation diagram showing the main organization of the important bifurcations. In particular, we can distinguish in Figure 17 regions where the positive equilibrium is unique and where it is not. We further notice that indeed, as shown in Appendix B, there is a Takens-Bogdanov bifurcation close to the point $T_{C}$. In there, we verify that the position of the Takens-Bogdanov point obtained numerically is indeed correct. In Figure 18, we show some representatives of the structure of the limit cycles in different regions of the parameter space. In particular, we verify the observation we made in … that, as soon as multiple positive equilibria exist, the limit cycles, either stable or unstable, limit a homoclinic orbit.
@@ Acknowledgements
Roberto Albarrán García was supported by a PhD scholarship from CONACYT, México.

### v2

@@ 6.3. Exit chart
Refer to caption $\varepsilon_{3}$$\varepsilon_{3}$$v_{3}$$v_{3}$$r_{3}$$r_{3}$$\ell_{3}$$\ell_{3}$$\mathcal{M}_{3}$$\Delta_{3}^{\textup{in}}$$\Delta_{3}^{\textup{out}}$
Figure 11. Dynamics of (35) in the chart $K_{3}=\{\bar{u}=1\}$ for $A-M+AM-\frac{1}{Q}<0$ on the left and $A-M+AM-\frac{1}{Q}>0$ on the right. The flow in blue occurs in the $(\varepsilon_{3},v_{3})$-plane, while the green one in the $(r_{3},v_{3})$-plane. The black orbits are a sketch of a sample of orbits flowing from $\Delta_{3}^{\textup{in}}$ and arriving at $\Delta_{3}^{\textup{out}}$. The red surface is the center manifold $\mathcal{P}_{3}$. Notice that since $\ell_{3}$ is simply the critical manifold $\mathcal{M}_{0}^{1}$ written in the local coordinates of this chart, the orbits on $\mathcal{P}_{3}$ can be identified with the slow manifolds that perturb from $\mathcal{M}_{0}^{1}$. Although we do not show $\Delta_{c}^{\textup{in}}$, as it is an exponentially thin strip, it is evident from the analysis presented that some of the orbits near $\mathcal{P}_{3}$ arrive at $\Delta_{3}^{\textup{out}}$.
@@ 6.4. Bottom chart
$u=r_{4}u_{4},\quad v=-r_{4},\quad\varepsilon=r_{4}\varepsilon_{4},$ (58)
@@
$\begin{split}\frac{\textup{d}r_{4}}{\textup{d}t}&=r_{4}\varepsilon_{4}(Q+u_{1})(A^{2}M+\mathcal{O}(r_{4})),\\ \frac{\textup{d}u_{4}}{\textup{d}t}&=-AMQ(1+(A+M-AM)u_{4})u_{4}-A^{2}M(Q+u_{4})u_{4}\varepsilon_{4}+\mathcal{O}(r_{4}),\\ \frac{\textup{d}\varepsilon_{4}}{\textup{d}t}&=-\varepsilon_{4}^{2}(Q+u_{4})(A^{2}M+\mathcal{O}(r_{4})).\end{split}$ (59)
@@
Refer to caption $\varepsilon_{1}$$u_{4}$$r_{4}$
Figure 12. Dynamics of (59) near the origin. In contrast to the flow of (35), the center manifold in the plane $(v_{4},\varepsilon_{4})$ is not unique.
6.5. Blown-up dynamics at $T_{\mathcal{C}}$, relaxation oscillations, and transitory canard
@@
At the singular level, i.e. for $\varepsilon=0$, the map $\Pi_{0\to 2}$ is described by the concatenation of the corresponding orbits for $\bar{r}=0$ in the blow-up analysis performed in Sections 6.1-6.4, refer in particular to Figures 9, 10 and 11 and the analysis leading to them, together with a blow-down. For the map $\Pi_{1\to 2}$ one just needs to be careful to have the appropriate parameters such that the slow flow along $\mathcal{M}_{0}^{1,r}$ is directed toward $T_{\mathcal{C}}$ as in Figure 7 and that the parameter $Q$ is chosen so that there is a maximal canard at the fold point $P$. This is achieved for parameters exponentially close to $Q=Q_{H}$. Since we have already shown in Section 5 that the singular Hopf bifurcation at $P$ for $Q=Q_{H}$ may be supercritical, the appropriate choice of parameters guarantees that such a maximal canard is stable for $\varepsilon>0$ sufficiently small. ∎
@@
For $A-M+AM<\frac{1}{Q}$, one can show that the trajectories that cross $\Sigma_{0}$ transversely converge to the origin. Since this does not lead to oscillatory behavior, we do not detail this case anymore. However, it is interesting to observe that, in this case, the degenerate point $T_{\mathcal{C}}$ is a saddle but eventually attracts all orbits.
@@
Consider the system (2) and let $C=-AMQ$ and $A-M+AM>\frac{1}{Q}$.
@@
Figure 15. Simulations of (2) with $S=\varepsilon=0.05$, $A=\frac{1}{2}$, $M=-\frac{1}{10}$, and $C=-AMQ$. The equilibrium point $E_{1}$ is located close to the fold point $P$, and is indicated in blue. The parameter $Q$ is chosen as $Q=Q_{H}-\delta$ where $\delta$ is shown at the top of each figure. Notice, indeed, the canard explosion occurring for the chosen parameters. We highlight that some of the canard cycles pass through the degenerate point $T_{\mathcal{C}}$, indicated in red.
@@ 8. Numerical analysis and validation
For our simulations, we take the following setup of values of the parameters: $A=\frac{1}{2}$, $M=-\frac{1}{10}$, $\varepsilon=\frac{1}{20}$, while $(C,Q)$ are going to be bifurcation parameters. In Figure 17, we present a codimension $2$ bifurcation diagram showing the main organization of the important bifurcations. In particular, we can distinguish in Figure 17 regions where the positive equilibrium is unique and where it is not. We further notice that indeed, as shown in Appendix B, there is a Takens-Bogdanov bifurcation close to the point $T_{\mathcal{C}}$. In there, we verify that the position of the Takens-Bogdanov point obtained numerically is indeed correct. In Figure 18, we show some representatives of the structure of the limit cycles in different regions of the parameter space. In particular, we observe that as soon as there are multiple equilibria, recall Remark 2, the limit cycles corresponding to a canard explosion end up in a homoclinic orbit. The rigorous analysis of (16) when multiple equilibria exist is a topic of future research.
@@ Acknowledgements
Roberto Albarran García was partially supported by a CONAHCyT Mexico postgraduate fellowship No. 924692.